# Fully Kinetic Simulations of Proton-Beam-Driven Instabilities from Parker Solar Probe Observations

### v1

@@ 3 Results
All the energies in Figure 4 we plot the difference in each energy $\delta E=E(t)-E(t=0)$ and normalized by $U_{0}$, which is the total energy of the system at initialization, using the formula $\delta E/U_{0}\doteq\left(E(t)-E(t_{0})\right)/U_{0}$. The quasi-stationary phase is characterized by a transient of quiescence in which the total magnetic energy $E^{\mathrm{m}}\doteq\int B^{2}/(8\pi)~{}\mathrm{d}^{3}\boldsymbol{x}$ and the total kinetic energy of electrons and protons $E^{\mathrm{k}}=\sum_{s}\left(E^{\mathrm{th}}_{s}+E^{\mathrm{d}}_{s}\right)$ remain nearly constant over time, as shown in Figure 4(a). The kinetic thermal energy for the species $s$ is represented as $E^{\mathrm{th}}_{s}\doteq 1/2\int\mathrm{Tr}(\boldsymbol{P}_{s})~{}\mathrm{d}^{3}\boldsymbol{x}$, where $\mathrm{Tr}(\boldsymbol{P}_{s})$ is the trace of the pressure tensor. The drift (i.e. bulk) energy of the generic species $s$ is defined as $E^{\mathrm{d}}_{s}\doteq 1/2\int\rho_{s}V_{s}^{2}~{}\mathrm{d}^{3}\boldsymbol{x}$, where $\boldsymbol{V}_{s}\doteq 1/n_{s}\int\boldsymbol{v}f_{s}~{}\mathrm{d}^{3}\boldsymbol{v}$ represent the drift velocity of the $s$-th species and $f_{s}$ denotes the distribution function of the $s$-th species. The pressure tensor is denoted as $\boldsymbol{P}_{s}\doteq m_{s}\int(\boldsymbol{v}-\boldsymbol{V}_{s})(\boldsymbol{v}-\boldsymbol{V}_{s})f_{s}~{}\mathrm{d}^{3}\boldsymbol{v}$, where $\boldsymbol{v}$ denotes the velocity. In a reference system aligned with the direction $\hat{\boldsymbol{b}}\doteq\boldsymbol{B}/\lVert\boldsymbol{B}\rVert$ of the magnetic field, the perpendicular pressure is thus defined as $p_{\perp,s}\doteq\boldsymbol{P}_{s}:\left(\mathbb{1}-\hat{\boldsymbol{b}}\hat{\boldsymbol{b}}\right)/2$, and the parallel pressure is $p_{\parallel,s}\doteq\boldsymbol{P}_{s}:\left(\hat{\boldsymbol{b}}\hat{\boldsymbol{b}}\right)$. The thermal energy for the general species $s$, in the magnetic field-aligned reference frame, is calculated in the perpendicular direction as $E^{\mathrm{th}}_{\perp,s}\doteq 1/(2k_{\mathrm{B}})\int p_{\perp,s}\,\mathrm{d}^{3}\boldsymbol{x}$ and in the parallel direction as $E^{\mathrm{th}}_{\parallel,s}\doteq 1/(2k_{\mathrm{B}})\int p_{\parallel,s}\,\mathrm{d}^{3}\boldsymbol{x}$. The components of kinetic energy in Figure 4(b) are: the beam parallel and perpendicular thermal energy, respectively $\delta E^{\mathrm{th}}_{\parallel,\mathrm{b}}/U_{0}$ and $\delta E^{\mathrm{th}}_{\perp,\mathrm{b}}/U_{0}$, core parallel and perpendicular thermal energy, respectively $\delta E^{\mathrm{th}}_{\parallel,\mathrm{c}}/U_{0}$ and $\delta E^{\mathrm{th}}_{\perp,\mathrm{c}}/U_{0}$, and the drift energy of both the core and the beam $\delta E^{\mathrm{d}}_{\mathrm{c}}/U_{0}$ and $\delta E^{\mathrm{d}}_{\mathrm{b}}/U_{0}$.
During the non-linear phase, the magnetic energy is converted back into kinetic energy, as illustrated in Figure 4(a). In Figure 4(b), it is evident that $\delta E^{\mathrm{th}}_{\perp,\mathrm{c}}/U_{0}$ and $\delta E^{\mathrm{th}}_{\parallel,\mathrm{c}}/U_{0}$ do not change significantly. Meanwhile, the $\delta E^{\mathrm{th}}_{\perp,\mathrm{b}}/U_{0}$ is steadily increasing, while $\delta E^{\mathrm{d}}_{\mathrm{b}}/U_{0}$ is consistently decreasing. During this phase, the bulk energy stored in the beam’s drift motion is converted into the thermal perpendicular energy of the beam. In Figure 4(c), the time evolution of the proton temperature ratio, defined as $T_{\perp,\mathrm{p}}/T_{\parallel,\mathrm{p}}=p_{\perp,\mathrm{p}}/p_{\parallel,\mathrm{p}}$ and integrated across the spatial domain, is presented. The perpendicular and parallel pressures for the total population of protons are calculated from from the pressure tensor $\boldsymbol{P}_{\mathrm{p}}\doteq m_{\mathrm{p}}\int(\boldsymbol{v}-\boldsymbol{V}_{\mathrm{cm},\mathrm{p}})(\boldsymbol{v}-\boldsymbol{V}_{\mathrm{cm},\mathrm{p}})f_{\mathrm{p}}\,\mathrm{d}^{3}\boldsymbol{v}$, where the velocity of the center of mass of the proton population is $\boldsymbol{V}_{\mathrm{cm},\mathrm{p}}\doteq\left(n_{\mathrm{c}}\boldsymbol{V}_{\mathrm{c}}+n_{\mathrm{b}}\boldsymbol{V}_{\mathrm{b}}\right)/\left(n_{\mathrm{c}}+n_{\mathrm{b}}\right)$. From these results, we observe that the pronounced outcome of the considered instability is the increase in proton temperature ratio, redistributing the energy from the parallel to the perpendicular direction. In conclusion, studying the energy exchanges of the system in Figure 4(a)-(c) reveals how energy is redistributed inside the system. The two primary sources of free energy are the beam temperature ratio and the relative drifting speed of the core and beam. Therefore, based on this analysis, we can conclude that the system exhibits an instability which we term as “beam-firehose instability” because it is powered not only by thermal anisotropy but also by the drifting population, resembling a beam-plasma type of instability.
@@
Figure 5: Evolution of the proton VDF $f_{\mathrm{p}}(v_{x},v_{y},t)$ (left column) and its parallel cut $f_{\mathrm{p}}(v_{x},v_{y}=0,t)$ (right column). Panels (a)–(b) represent the quasi-stationary phase of the proton-beam instability instability. Panel (a) shows the initial distribution of protons $f_{\mathrm{p}}(v_{x},v_{y},t)$, whereas panel (b) shows the horizontal cut $f_{\mathrm{p}}(v_{x},v_{y}=0,t)$ (with the core and beam sub-populations indicated by different lines), at $t/\Omega_{\mathrm{p}}^{-1}=0$. Panels (c)–(d) represent the phase space of protons during the linear phase at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$. The proton VDF during the non-linear phase at $t/\Omega_{\mathrm{p}}^{-1}\approx 436$ is shown in panels (e) and (f). The resonant velocity of the system, $v_{\mathrm{res}}=(\omega_{\mathrm{r}}+\Omega_{\mathrm{p}})/k_{\parallel}$, is shown by the blue line in panel (f).
@@
Figure 5 shows the time evolution of the core, beam, and total proton VDF, denoted as $f_{\mathrm{p}}(v_{x},v_{y},t)$, in the left column, as well as a one-dimensional cut $f_{\mathrm{p}}(v_{x},v_{y}=0,t)$, in the right column, to provide further evidence of the wave-particle resonant interaction. Panels (a) and (b) in Figure 5 represent the initial conditions of the simulation, corresponding to the beginning of the quasi-stationary phase of the instability. Subsequently, panels (c) and (d) in Figure 5 at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$ illustrate the linear phase, during which the beam population undergoes resonant interaction with the excited fast magnetosonic waves. Figure 5(e)-(f) represents the proton VDF during the non-linear phase at $t/\Omega_{\mathrm{p}}^{-1}\approx 436$. A hint of the “hammerhead-like” structure (extensions of the beam towards the values of higher $v_{y}$), initially observed forming during the linear phase in Figure 5(c) and (d), becomes more pronounced during the non-linear phase in Figure 5(e)–(f). In Figure 5(d) and (f), we have shown a vertical line indicating the resonant velocity $v_{\mathrm{res}}\doteq(\omega_{\mathrm{r}}-n\Omega_{\mathrm{p}})/k_{\parallel}$, where $n$ is any integer, for the fast-magnetosonic wave quantities: $k_{\parallel}\approx 0.5$, $\Omega_{\mathrm{p}}\approx 0.0029$, $\omega_{\mathrm{r}}\approx 0.0015$. Cyclotron resonance with ${n}=-1$, which corresponds to the backward-propagating Alfvén wave and the forward-propagating fast-magnetosonic mode (Klein et al., 2021), has been assumed. Comparing this to the distribution in Figure 5(b), we observe a decrease in the number of particles for velocities $v>v_{\mathrm{res}}$ which identifies a phase space depletion region. Conversely, we observe an increase in the number of particles for velocities $v<v_{\mathrm{res}}$ creating a phase space pile-up region.
@@
Figure 6: Proton VDF $f_{\mathrm{p}}(v_{x},v_{y},t)$ at time $t/\Omega_{\mathrm{p}}^{-1}\approx 436$ normalised by its value at the initial time. The blue line shows the parallel resonant velocity $v_{\mathrm{r}}$. The black dashed circle centered at $(v_{x},v_{y})=(v_{\mathrm{ph}},0)$, with $v_{\mathrm{ph}}=\omega_{\mathrm{r}}/k_{\parallel}$ the parallel phase velocity of the waves, is the diffusion path due to the resonant interaction.
@@
Here, we will investigate the nature of proton diffusion through wave-particle resonant interaction. In Figure 6, the proton VDF is shown at time $t/\Omega_{\mathrm{p}}^{-1}=436$, normalised by its value at the initial time. The normalization allows us to see small changes in the VDF more easily. In the white regions, no significant particle diffusion occurs, while in the blue regions, there are fewer particles than in the initial state, and the red regions are where particles have accumulated. The blue line here indicates the parallel resonant velocity, while the black circle, $(v_{x}-v_{\mathrm{ph}})^{2}+v_{y}^{2}=1$, represents an iso-line defining the path in velocity phase space where particle energy remains constant. Here, $v_{\mathrm{ph}}=\omega_{\mathrm{r}}/k_{\parallel}$ denotes the parallel phase velocity of the waves. Particles diffuse in phase space along trajectories tangent to the iso-line of constant energy. Resonant particles experience a gain in kinetic energy, manifested as an increase in $v_{x}^{2}+v_{y}^{2}$. This acquired energy is extracted from the waves at $\omega_{\mathrm{res}}$, resulting in their damping. As illustrated in Figure 6, a VDF depletion region (blue) is observed for velocities $v>v_{\mathrm{res}}$, while a phase space pile-up region occurs for $v<v_{\mathrm{res}}$.
@@ 4 Discussion and Conclusions
Based on our analysis of the temporal evolution of the energy, we can summarise a few key points. The proton core sub-population does not contribute significantly to the dynamics of the system, which is primarily dominated by the beam sub-population. The thermal energy of the beam protons is transferred from the parallel to the perpendicular direction. The parallel thermal energy of the proton beam increases at the expense of the proton beam drift energy, which constitutes the main source of free energy. During the nonlinear phase, the beam drift energy is contributing to the increase in perpendicular thermal energy of the beam. The notable result of the wave-particle interaction is an increase in proton temperature ratio, redistributing energy from the parallel to the perpendicular direction.
@@
In our simulated proton VDFs, starting from the linear phase, we note a reduction in the number of particles for velocities exceeding the resonant velocity $v_{\mathrm{res}}$, indicating a phase space depletion region. Conversely, there is an increase in the particle count for velocities below $v_{\mathrm{res}}$, manifesting as a phase space pile-up region. Our simulation indicates that the fast magnetosonic waves are inducing resonance in the proton beam population. Consequently, these protons experience diffusion in phase space, leading to perpendicular heating effects as they move into less densely populated regions, resulting in larger perpendicular velocity (Verscharen et al., 2019a). The diffusion process initiates in the linear phase and continues into the nonlinear phase, where it saturates, gradually diminishing in intensity.
@@
Cyclotron resonance plays a significant role in shaping magnetic field spectra and distribution functions observed in solar wind (Goldstein et al., 1994; Leamon et al., 1998). The nonlinear evolution of the fast magnetosonic mode is significantly influenced by cyclotron resonance, which plays a crucial role in inducing perpendicular proton heating and redistributing energy from the parallel to the perpendicular direction. This resonance process effectively mitigates the free energy associated with the parallel anisotropy, constituting a new channel of regulation of the solar wind temperature anisotropy. We demonstrated that wave-particle resonant interaction can modify the proton beam VDF producing a hint of a “hammerhead” distribution. This supports the hypothesis regarding the origin of the proton hammerhead (i.e. strongly asymmetric beam) VDF proposed by Verniero et al. (2022). The width of the hammerhead-like feature we see in the proton VDFs, induced by the resonant interaction between fast magnetosonic waves and protons, is smaller with respect to observations Verniero et al. (2022). We speculate that the presence of turbulence, which is ubiquitous in the solar wind, could provide an additional source of energy to be dissipated via the interaction with the proton beam, eventually producing a more pronounced hammerhead. It is known that turbulence drives ion heating (Servidio et al., 2012; Matthaeus et al., 2020; Arrò et al., 2022), inducing anisotropic deformations on Maxwellian particle distributions (Servidio et al., 2015; Arrò et al., 2023). However, the interplay between turbulence and non-Maxwellian distributions, like those studied in the present work, has been poorly investigated and could significantly impact dissipation. This point will be addressed in future works.
@@
The paper by Ofman et al. (2022) presents a comparison between 2D and 3D hybrid simulations of ion beam instability in solar wind plasma, based on observations from the PSP near perihelia. The study demonstrates that the 3D simulation does not exhibit additional effects compared to the 2D simulation, indicating that the third dimension is not necessary to study proton-beam-driven instabilities. However, solar wind protons may be more accurately described by a kappa distribution function instead of a Maxwellian distribution (Livadiotis & McComas, 2013; Nicolaou et al., 2020). A natural progression of this work could involve extending the simulation with a study of this system modeling the proton beam using a kappa distribution. However, these aspects are beyond the scope of this paper and are left for future studies.

### v2

@@ 3 Results
In Figure 4(a)-(c) we plot the difference in each energy $\delta E=E(t)-E(t=0)$ normalized by $U_{0}$, which is the total energy of the system at initialization, using the formula $\delta E/U_{0}\doteq\left(E(t)-E(t_{0})\right)/U_{0}$. The quasi-stationary phase is characterized by a transient in which the total magnetic energy $E^{\mathrm{m}}\doteq\int B^{2}/(8\pi)~{}\mathrm{d}^{3}\boldsymbol{x}$ and the total kinetic energy of electrons and protons $E^{\mathrm{k}}=\sum_{s=\mathrm{e},\mathrm{p}}\left(E^{\mathrm{th}}_{s}+E^{\mathrm{d}}_{s}\right)$ remain nearly constant over time, as shown in Figure 4(a). The kinetic thermal energy for species $s$ is defined as $E^{\mathrm{th}}_{s}\doteq(1/2)\int\mathrm{Tr}(\boldsymbol{P}_{s})~{}\mathrm{d}^{3}\boldsymbol{x}$, where $\mathrm{Tr}(\boldsymbol{P}_{s})$ is the trace of the pressure tensor for species $s$. The drift (i.e. bulk) energy of the generic species $s$ is defined as $E^{\mathrm{d}}_{s}\doteq(1/2)\int\rho_{s}V_{s}^{2}~{}\mathrm{d}^{3}\boldsymbol{x}$, where $\boldsymbol{V}_{s}\doteq(1/n_{s})\int\boldsymbol{v}f_{s}~{}\mathrm{d}^{3}\boldsymbol{v}$ represents the drift velocity of the $s$-th species and $f_{s}$ denotes the distribution function of the $s$-th species. The pressure tensor is given by $\boldsymbol{P}_{s}\doteq m_{s}\int(\boldsymbol{v}-\boldsymbol{V}_{s})(\boldsymbol{v}-\boldsymbol{V}_{s})f_{s}~{}\mathrm{d}^{3}\boldsymbol{v}$, where $\boldsymbol{v}$ denotes the particle velocity. In a reference system aligned with the direction $\hat{\boldsymbol{b}}\doteq\boldsymbol{B}/\lVert\boldsymbol{B}\rVert$ of the magnetic field, the perpendicular pressure is thus defined as $p_{\perp,s}\doteq\boldsymbol{P}_{s}:(\mathbb{1}-\hat{\boldsymbol{b}}\hat{\boldsymbol{b}})/2$, and the parallel pressure is $p_{\parallel,s}\doteq\boldsymbol{P}_{s}:\hat{\boldsymbol{b}}\hat{\boldsymbol{b}}$. The thermal energy for the generic species $s$, in the magnetic field-aligned reference frame, is calculated in the perpendicular direction as $E^{\mathrm{th}}_{\perp,s}\doteq(1/(2k_{\mathrm{B}}))\int p_{\perp,s}\,\mathrm{d}^{3}\boldsymbol{x}$ and in the parallel direction as $E^{\mathrm{th}}_{\parallel,s}\doteq(1/(2k_{\mathrm{B}}))\int p_{\parallel,s}\,\mathrm{d}^{3}\boldsymbol{x}$. The components of kinetic energy in Figure 4(b) are: the beam parallel and perpendicular thermal energy, respectively $\delta E^{\mathrm{th}}_{\parallel,\mathrm{b}}/U_{0}$ and $\delta E^{\mathrm{th}}_{\perp,\mathrm{b}}/U_{0}$, core parallel and perpendicular thermal energy, respectively $\delta E^{\mathrm{th}}_{\parallel,\mathrm{c}}/U_{0}$ and $\delta E^{\mathrm{th}}_{\perp,\mathrm{c}}/U_{0}$, and the drift energy of both the core and the beam $\delta E^{\mathrm{d}}_{\mathrm{c}}/U_{0}$ and $\delta E^{\mathrm{d}}_{\mathrm{b}}/U_{0}$.
During the non-linear phase, the magnetic energy is converted back into kinetic energy, as illustrated in Figure 4(a). As typically observed after instability saturation, the driver behind the instability onset is removed by the instability itself. In Figure 4(b), it is evident that $\delta E^{\mathrm{th}}_{\perp,\mathrm{c}}/U_{0}$ and $\delta E^{\mathrm{th}}_{\parallel,\mathrm{c}}/U_{0}$ do not change significantly. Meanwhile, $\delta E^{\mathrm{th}}_{\perp,\mathrm{b}}/U_{0}$ is steadily increasing, while $\delta E^{\mathrm{d}}_{\mathrm{b}}/U_{0}$ is consistently decreasing. During this phase, the bulk energy stored in the beam’s drift motion is converted into thermal perpendicular energy of the beam. A key parameter in proton beam instabilities is the core-to-beam drift velocity, which determines the growth rate of the instability and the types of waves produced (Shaaban et al., 2020). The electron population, as shown in Figure 4(c), does not significantly contribute to the system’s dynamics. The electron bulk energy remains nearly constant around zero, with a slight increase during the quasi-linear phase. The thermal energy, instead, monotonically increases throughout the entire time of the simulation. However, the magnitudes of electron drift and thermal energy are notably smaller compared to those of the protons (both core and beam). Therefore, we conclude that there is no interplay between protons and electrons in our case.
In Figure 4(d)-(f), we present the time evolution of the proton temperature anisotropy for core, beam and the total proton populations. In the latter case, we consider core plus beam as a single proton population, and we calculate the moments of this single population rather than the moments for core and beam separately. The perpendicular and parallel pressures for the total population of protons are calculated from the pressure tensor $\boldsymbol{P}_{\mathrm{p}}\doteq m_{\mathrm{p}}\int(\boldsymbol{v}-\boldsymbol{V}_{\mathrm{cm},\mathrm{p}})(\boldsymbol{v}-\boldsymbol{V}_{\mathrm{cm},\mathrm{p}})f_{\mathrm{p}}\,\mathrm{d}^{3}\boldsymbol{v}$, where the velocity of the center of mass of the proton population is $\boldsymbol{V}_{\mathrm{cm},\mathrm{p}}\doteq\left(n_{\mathrm{c}}\boldsymbol{V}_{\mathrm{c}}+n_{\mathrm{b}}\boldsymbol{V}_{\mathrm{b}}\right)/\left(n_{\mathrm{c}}+n_{\mathrm{b}}\right)$. From these results, shown in panels (d)-(f), we infer that the pronounced outcome of the considered instability is the increase in proton temperature anisotropy in the quasi-linear phase, for both the core (Figure 4(d)) and the beam (Figure 4(d)) proton populations, and consequently this is reflecting on the total proton population (Figure 4(f)). Therefore, the energy is redistributed from the parallel to the perpendicular direction.
In conclusion, studying the energy exchanges of the system in Figure 4(a)-(f) reveals how energy is redistributed inside the system. The two primary sources of free energy are the beam temperature anisotropy and the relative drift speed of the core and beam. Therefore, based on this analysis, we can conclude that the system evolves through a proton-beam instability, also known as ion-ion instability, which is powered not only by thermal anisotropy but also by the drifting population, resembling a beam-plasma type of instability.
@@
Figure 5: Evolution of the proton VDF $f_{\mathrm{p}}(v_{x},v_{y},t)$ (left column) and its parallel cut $f_{\mathrm{p}}(v_{x},v_{y}=0,t)$ (right column). Panels (a)–(b) represent the quasi-stationary phase of the proton-beam instability instability. Panel (a) shows the initial distribution of protons $f_{\mathrm{p}}(v_{x},v_{y},t)$, whereas panel (b) shows the horizontal cut $f_{\mathrm{p}}(v_{x},v_{y}=0,t)$ (with the core and beam sub-populations indicated by different lines), at $t/\Omega_{\mathrm{p}}^{-1}=0$. Panels (c)–(d) represent the phase space of protons during the quasi-linear phase at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$. The proton VDF during the non-linear phase at $t/\Omega_{\mathrm{p}}^{-1}\approx 436$ is shown in panels (e) and (f). The resonant velocity of the system, $v_{\mathrm{res}}=(\omega_{\mathrm{r}}+\Omega_{\mathrm{p}})/k_{\parallel}$, is shown by the blue line in panel (f).
@@
Figure 5 shows the time evolution of the core, beam, and total proton VDF, denoted as $f_{\mathrm{p}}(v_{x},v_{y},t)$, in the left column, as well as a one-dimensional cut $f_{\mathrm{p}}(v_{x},v_{y}=0,t)$, in the right column, to provide further evidence of the wave-particle resonant interaction. Panels (a) and (b) in Figure 5 represent the initial conditions of the simulation, corresponding to the beginning of the quasi-stationary phase of the instability. Subsequently, panels (c) and (d) in Figure 5 at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$ illustrate the quasi-linear phase, during which the beam population undergoes resonant interaction with the excited fast magnetosonic waves. Figure 5(e)-(f) represents the proton VDF during the non-linear phase at $t/\Omega_{\mathrm{p}}^{-1}\approx 436$. A hint of the anisotropic proton beam (i.e. extensions of the beam towards the values of higher $v_{y}$), initially observed forming during the quasi-linear phase in Figure 5(c) and (d), becomes more pronounced during the non-linear phase in Figure 5(e)–(f). In Figure 5(d) and (f), we have shown a vertical line indicating the resonant velocity $v_{\mathrm{res}}\doteq(\omega_{\mathrm{r}}-o\Omega_{\mathrm{p}})/k_{\parallel}$, where $o\in\mathbb{N}$, for the fast-magnetosonic wave quantities: $k_{\parallel}/d_{\mathrm{p}^{-1}}\approx 0.501$, $\Omega_{\mathrm{p}}/\omega_{\mathrm{p}}\approx 0.0029$, $\omega_{\mathrm{r}}/\omega_{\mathrm{p}}\approx 0.0015$. Cyclotron resonance with ${o}=-1$, which corresponds to the backward-propagating Alfvén wave and the forward-propagating fast-magnetosonic mode (Klein et al., 2021), has been assumed. Comparing this to the distribution in Figure 5(b), we observe a decrease in the number of particles for velocities $v>v_{\mathrm{res}}$ which identifies a phase space depletion region. Conversely, we observe an increase in the number of particles for velocities $v<v_{\mathrm{res}}$ creating a phase space pile-up region.
@@
Figure 6: Proton VDF $f_{\mathrm{p}}(v_{x},v_{y},t)$ at time $t/\Omega_{\mathrm{p}}^{-1}\approx 436$ normalized by its value at the initial time. The blue line shows the parallel resonant velocity $v_{\mathrm{r}}$. The black dashed circle centered at $(v_{x},v_{y})=(v_{\mathrm{ph}},0)$, with $v_{\mathrm{ph}}=\omega_{\mathrm{r}}/k_{\parallel}$ the parallel phase velocity of the waves, is the diffusion path due to the resonant interaction.
@@
Here, we will investigate the nature of proton diffusion through wave-particle resonant interaction. In Figure 6, the proton VDF is shown at time $t/\Omega_{\mathrm{p}}^{-1}=436$, normalized by its value at the initial time. The normalization allows us to highlight small changes in the VDF more easily. In the white regions, no significant particle diffusion occurs, while in the blue regions, there are fewer particles than in the initial state, and the red regions are where particles have accumulated. The blue vertical line here indicates the parallel resonant velocity, while the black circle, $(v_{x}-v_{\mathrm{ph}})^{2}+v_{y}^{2}=1$, represents an iso-line defining the path in velocity phase space where particle energy remains constant. Here, $v_{\mathrm{ph}}=\omega_{\mathrm{r}}/k_{\parallel}$ denotes the parallel phase velocity of the waves. Particles diffuse in phase space along trajectories tangent to the iso-line of constant energy. Resonant particles experience a gain in kinetic energy, manifested as an increase in $v_{x}^{2}+v_{y}^{2}$. This acquired energy is extracted from the waves at resonant frequency $\omega_{\mathrm{res}}$, resulting in their damping. As illustrated in Figure 6, a VDF depletion region (blue) is observed for velocities $v>v_{\mathrm{res}}$, while a phase space pile-up region occurs for $v<v_{\mathrm{res}}$.
@@ 4 Discussion and Conclusions
Based on our analysis of the temporal evolution of the energy, we can summarise a few key points. The proton core sub-population does not contribute significantly to the dynamics of the system, which is primarily dominated by the beam sub-population. The thermal energy of the beam protons is transferred from the parallel to the perpendicular direction. The parallel thermal energy of the proton beam increases at the expense of the proton beam drift energy, which constitutes the main source of free energy. During the non-linear phase, the beam drift energy is contributing to the increase in perpendicular thermal energy of the beam. The notable result of the wave-particle interaction is an increase in proton temperature anisotropy, redistributing energy from the parallel to the perpendicular direction.
@@
In our simulated proton VDFs, starting from the quasi-linear phase, we note a reduction in the number of particles for velocities exceeding the resonant velocity $v_{\mathrm{res}}$, indicating a phase space depletion region. Conversely, there is an increase in the particle count for velocities below $v_{\mathrm{res}}$, manifesting as a phase space pile-up region. Our simulation indicates that the fast magnetosonic waves are inducing resonance in the proton beam population. Consequently, these protons experience diffusion in phase space, leading to perpendicular heating effects as they move into less densely populated regions, resulting in larger perpendicular velocity (Verscharen et al., 2019a). The diffusion process initiates in the quasi-linear phase and continues into the non-linear phase, where it saturates, gradually, diminishing in intensity.
@@
Cyclotron resonance plays a significant role in shaping magnetic field spectra and distribution functions observed in the solar wind (Goldstein et al., 1994; Leamon et al., 1998). The nonlinear evolution of the fast magnetosonic mode is significantly influenced by cyclotron resonance, which plays a crucial role in inducing perpendicular proton heating and redistributing energy from the parallel to the perpendicular direction. This resonance process effectively mitigates the free energy associated with the parallel anisotropy, constituting a new channel of regulation of the solar wind temperature anisotropy. We demonstrated that wave-particle resonant interaction can modify the proton beam VDF, inducing perpendicular heating that reaches up to 12-15% of its original value. According to Verniero et al. 2022, the “hammerhead” features in the observed VDFs are the result of proton beam driven instabilities of the parallel FM/W instability. The strongly asymmetric proton beam exhibits a temperature anisotropy in the range of 2 to 8. We speculate that our simulation could produce a more pronounced proton beam anisotropy by considering the presence of turbulence, which is ubiquitous in the solar wind. This turbulence could provide an additional energy source, dissipated through interaction with the proton beam, thereby enhancing the anisotropy. It is known that turbulence drives ion heating (Servidio et al., 2012; Matthaeus et al., 2020; Arró et al., 2022), inducing anisotropic deformations on Maxwellian particle distributions (Servidio et al., 2015; Arrò et al., 2023, 2024). However, the interplay between turbulence and non-Maxwellian distributions, like those studied in the present work, has been poorly investigated and could significantly impact dissipation. This point will be addressed in future works.
@@
The work by Ofman et al. 2022 presents a comparison between 2D and 3D hybrid simulations of ion beam instability in solar wind plasma, based on observations from the PSP near perihelia. That study demonstrated that 3D simulations do not exhibit additional effects compared to 2D simulations, indicating that three-dimensional effects do not significantly affect proton-beam-driven instabilities. However, solar wind protons may be more accurately described by a kappa distribution function instead of a Maxwellian distribution (Pierrard & Lazar, 2010; Livadiotis & McComas, 2013; Nicolaou et al., 2020). A natural progression of this work could involve extending the simulation with a study of this system modeling the proton beam using a kappa distribution. However, these aspects are beyond the scope of this paper and are left for future studies.